Status Report

Hubble Space Telescope Daily Report #2741 11/02/00

By SpaceRef Editor
November 2, 2000
Filed under

PERIOD COVERED:  0000Z (UTC) 11/01/00 – 0000Z (UTC) 11/02/00
Daily Status Report as of 307/0000Z
    1.1 Completed Three Sets of WF/PC-2 8826 (Cycle 9 Supplemental Darks pt 1)
        The WF/PC-2 was used to perform a T dark calibration program that obtains three dark frames every day in order to provide data for monitoring and characterizing the evolution of hot pixels.  There were no reported anomalies.
    1.2 Completed Two Sets of STIS/CCD 8837 (CCD Dark Monitor-Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks.  The proposal completed with no reported problems.
    1.3 Completed Three Sets of STIS/CCD 8631 (Bright Quasar Close Lensing Search II)
        The Space Telescope Imaging Spectrograph (CCD) was used to expand the Cycle 8 second generation HST snapshot survey of bright quasars, optimized to find lenses with component image separations < 1".  The observations completed nominally.
    1.4 Completed WF/PC-2 8632 (A UV Atlas of Nearby Galaxies)
        The WF/PC-2 was used to perform a snapshot survey of local galaxies at UV wavelengths with the F300W filter.  The aim of the project is to build a reference UV Atlas of normal galaxies, whose optical images are well known, with the highest possible degree of information, covering all the morphological types and luminosity classes.  The proposal completed normally.
    1.5 Completed Two Sets of STIS/CCD 8562 (Probing the Large Scale Structure: Cosmic Shear Observations)
        The Space Telescope Imaging Spectrograph (CCD) was used to probe the distortion of light bundles from distant galaxies, looking at the statistical properties of the intervening inhomogeneous {dark} matter distribution.  The proposal completed nominally.
    1.6 Completed STIS/CCD 8838 (Bias Monitor – Part 1)
        The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.  The proposal completed nominally.
    1.7 Completed STIS/CCD 8664 (Structural Measurement of Globular Clusters in M31 and NGC 5128: Stalking the Fundamental Plane)
        The Space Telescope Imaging Spectrograph (CCD) was used to observe in the snapshot mode to image a wide selection of individual globular clusters in two other large galaxies {M31 and NGC 5128} for measurement of their structural parameters {r_c, c, central surface brightness}.  We will use these to compute their binding energies and define the FP in these two galaxies. Comparison with the Milky Way will then give us powerful new information on just how “universal” the cluster formation process was in the early protogalaxies.  The observations completed with no reported problems.
    1.8 Completed STIS/CCD 7462 (Velocity Structure in High Redshift Galaxies)
        The Space Telescope Imaging Spectrograph (CCD) was used to observe a recently-discovered optical Einstein ring.  The presence of a strong Ly_Alpha emission line in the spectrum of the ring allows one to map the position of the emission line peak around the ring, thereby measuring velocity differences in the extended source object, a star-forming galaxy at z=3.6.  The observations completed nominally, with no reported problems.
    1.9 Completed Two Sets of WF/PC-2 8805 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation)
        The WF/PC-2 was used to perform a generic target version of the WFPC2 Archival Pure Parallel program.  The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group.  The proposal completed with no reported problems.
    1.10 Completed WF/PC-2 8822 (Decontaminations and Associated Observations Pt. 1/4)
        The Wf/PC-2 was used for the monthly decons.  Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, intflats, kspots, & darks}, UV throughput check, VISFLAT sweep, and internal UV flat check.  There were no reported problems.
    1.11 Completed WF/PC-2/STIS/CCD 8559 (The Role of Dark Matter in Cluster Formation and Galaxy Evolution)
        The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to observe the outer regions of massive clusters that represent transitional areas of great cosmological importance where field galaxies encounter the steep potential wells of dark matter and baryonic hot gas.  Little is known about either the dark matter profile at large radii or the morphological properties of infalling galaxies at those redshifts where strong evolution is observed in the cluster cores.  The proposal completed normally, with no reported problems
The following scheduled proposals were not executed due to the WF/PC-2 safing:
WF/PC-2 8586 (Observations of Astrophysically Important Visual Binaries) WF/PC-2 8826 (Cycle 9 Supplemental Darks pt 1)
    2.1 Guide Star Acquisitions:
      Scheduled Acquisitions:  4
Successful:  4
  Scheduled Re-acquisitions: 12
Successful: 12
    2.2 FHST Updates:
Scheduled: 10
Successful: 10
    2.3 Operations Notes:
        Using ROP SR-1A, the SSR EDAC error counter was cleared twice.
        SI C&DH errors were cleared at 306/2253Z and at 307/0255Z per ROP NS-5.
        The WF/PC-2 was successfully safed at 306/223507Z with real-time commanding contained in an operations request.  That decision was made in response to newly observed anomalous behavior of the shutter
mechanism.  The WF/PC-2 Anomaly Review Board has identified a modification to the WF/PC-2 slight software that can compensate for the erroneous sensing of the shutter blade position.  The timing associated with the blade position sensing can be extended to significantly increase the probability of capturing the true position of the shutter blade.  The flight software patch is scheduled to be installed on day 313 (November 8).
        Engineering status buffer limits were adjusted at 307/0205 as directed by ROP DF-18A.
        Continuation of normal science observations and calibrations.

SpaceRef staff editor.